# PromptIntern: Saving Inference Costs by Internalizing Recurrent Prompt during Large Language Model Fine-tuning

### v1

@@ Abstract
Large language models (LLMs) have played a fundamental role in various natural language processing tasks with powerful prompt techniques. However, in real-world applications, there are often similar prompt components for repeated queries, which causes significant computational burdens during inference. Existing prompt compression and direct fine-tuning methods aim to tackle these challenges, yet they frequently struggle to strike an optimal balance between cost-efficiency and performance effectiveness, especially in complex tasks such as NL2Code. In this paper, we propose a novel method namely PromptIntern to internalize the prompt knowledge into model parameters via progressive fine-tuning. Our method enables LLMs to emulate the human learning process for a new task, where detailed templates and examples in a prompt are gradually internalized and phased out progressively as the model grows accustomed to the task. Extensive experiments demonstrate that our method reduces inference tokens over 90%, speedups inference by 4.2 times, and saves 88.3% monetary cost.
PromptIntern: Saving Inference Costs by Internalizing Recurrent Prompt during Large Language Model Fine-tuning
@@ 1 Introduction
Large language models (LLMs) have become pivotal in numerous natural language processing (NLP) applications, such as natural language generation Dong et al. (2019), reasoning  Zhu et al. (2023); Sui et al. (2023), and code generation Luo et al. (2023); He et al. (2024); Rozière et al. (2024). To enhance the predictive accuracy of LLMs in domain-specific tasks, recent techniques in fine-tuning, such as parameter-efficient fine-tuning (PEFT) He et al. (2021); Hu et al. (2021); Lester et al. (2021), have been developed for pre-trained models to excel in specific tasks by adjusting their parameters to better align with targeted datasets Hu et al. (2021). Many of these fine-tuning approaches typically adopt prompts that are optimized and integrated with detailed instructions, examples, and retrieved documents through prompt engineering techniques such as chain-of-thought Wei et al. (2022), few-shot prompting Brown et al. (2020), and retrieval-augmented generation Lewis et al. (2020). Cheng et al. (2023).
@@
Although these advancements enhance the capabilities of LLMs during fine-tuning, they also present new challenges: Prompt engineering often necessitates longer prompts, and directly integrating lengthy prompts into training process further increases computational costs during inference VM et al. (2024). This increase in cost precludes LLMs in many cost-sensitive scenarios where computational resources are constrained. Several prompt compression methods Li et al. (2023); Jiang et al. (2023a); Pan et al. (2024) have been proposed to reduce text redundancy. They design various prompt compression systems and strive to preserve maximum information between original and compressed prompts. While these methods ensure the retention of original prompt information, they mainly focus on the prompt perplexity and overlook the adaption of target LLMs during compression Pan et al. (2024). For challenging tasks that require model fine-tuning Mosbach et al. (2023), these approaches struggle to establish connections between compressed tokens and dynamically adjusted model parameters. Consequently, naively adopting these compression methods often leads to significant performance degradation, as relevant information may be inadvertently removed or distorted during the compression process.
In this paper, we propose a novel prompt internalization approach, namely ”PromptIntern”, that internalizes the prompt input during model fine-tuning and enables efficient inference. Unlike prompt compression that removes tokens based solely on prompt information entropy, we aim to transfer various types of prompt knowledge into updated model parameters, thereby directly enhancing LLMs understanding. Our idea is motivated by the human learning process as illustrated in Figure 1: During on-boarding, human interns need detailed instructions, examples, and documents to learn new tasks. As they internalize these materials and become familiar with their duties, they master the necessary skills and no longer require extra guidance. Similarly, when specific prompt information (e.g. task constraints, output formats/styles) is repeatedly exposed to a LLM during fine-tuning, the model can gradually internalize the knowledge into its updated parameters. And such repeated information can be progressively eliminated from prompt inputs since it becomes unnecessary for inference of the master LLM.
@@
  • We propose PromptIntern, a novel prompt internalization method that aims to internalize repetitive prompt knowledge into the model’s parameters, achieving extreme inference efficiency while maintaining high performance.
  • We devise detailed prompt internalization strategies for template compression and example absorption along with a tailored progressive fine-tuning pipeline.
  • We conducted extensive experiments with detailed analyses on challenging NL2Code tasks. The experimental results show that our approach reduces token usage over 90%, speeds up inference time by 4.2 times, and achieves 88.3% cost savings across a broad spectrum of LLMs.
@@ 2 Related Work
Prompt compression
rephrases original prompts more concisely and is classified into task-aware and task-agnostic approaches. Task-aware approaches, like LongLLMLingua Jiang et al. (2023b), adopt a question-aware coarse-to-fine strategy, adapting information based on the query, while methods like soft prompts Wingate et al. (2022); Liu et al. (2022); Mu et al. (2024) use learnable tokens to condense prompts. Conversely, task-agnostic methods utilize metrics such as information entropy to eliminate redundant prompt information, with systems like LLMLingua Jiang et al. (2023a); Li et al. (2023) estimating token importance using a smaller model. Despite the demonstrated effectiveness of these methods, producing compressed text that can generalize across different tasks and be effectively integrated into training scenarios remains a challenge.
Model fine-tuning
adapts pre-trained LLMs to specific tasks by modifying parameters. Based on the assumption that fine-tuning adds less new information to the model pre-trained on large internet-scale datasets, Parameter-Efficient Fine-Tuning (PEFT) methods aim to curtail the costs of tuning large models by adjusting a subset of parameters. Existing PEFT methods can be broadly categorized into three main approaches: 1) Adapter-based methods Houlsby et al. (2019); He et al. (2021): Introduce trainable modules within a static "backbone" network, offering flexibility but potentially increasing model size. 2) Prompt-based methods Lester and Constant ; Razdaibiedina et al. (2023); Nashid et al. (2023): Employ trainable "soft tokens" at input sequence start, requiring effective prompt design per task. 3) Low-rank adaptation methods Hu et al. (2021); Dettmers et al. (2024); Liu et al. (2024): Use low-rank matrices to approximate required weight adjustments, avoiding additional inference burden and often delivering strong performance.
@@ 3 Problem Formulation
In this paper, we define a input prompt as $x=(x^{tmp},x^{egs},x^{que})$, where each input prompt $x$ is considered as a tuple of three components: $x^{tmp}$ as the template such as fixed instructions, API docs, etc., $x^{egs}$ as the examples, and $x^{que}$ as the query. Typically, $x^{tmp}$ and $x^{egs}$ are relatively fixed and lengthy but essential for complex tasks. Let $f_{\theta}(\cdot)$ denotes the neural network function of a LLM model, typically transformer Vaswani et al. (2017), parameterized by $\theta$. The generated output by LLM can be represented as $f_{\theta}(x)$.
@@ 4 Methodology
In this section, we introduce the detailed procedures of PromptIntern. We first discuss the template compression that is designed to compress the entire fixed template part inside an input prompt. Then we describe the example absorption on how to effectively absorb demonstration examples into model parameters. Finally, we introduce a tailored training strategy for PromptIntern. The overall framework is shown on Figure 2.
@@ 4.1 Template Compression
We first introduce template compression, which is designed to compress the common template information existed across training instances. The motivation of the template compression stems from the following aspects: 1) Redundancy. The instruction is repetitive across prompts for a given task, often containing unnecessary tokens that do not contribute to the language model’s understanding, posing significant memory and computational burdens when the instruction is lengthy; and 2) Noise. Excessively long prompts may incorporate extraneous elements—either irrelevant or misleading information—that serves as noise and can adversely affect the model’s generation.
@@
where $C$ is a specific template compressor, $\tilde{x}^{tmp}$ is the compressed template, and $\tau^{tmp}$ is the template compression rate as defined in Jiang et al. (2023a), varying at differnt training interations. We then adopt a predetermined schedule $\mathcal{S}^{tmp}(t)$ to progressively reduce and internalize the prompt template information during the $t$-th training iteration. Specifically, for a total of $T$ training iterations, we initially set $\tau^{tmp}$ to 1 at $\mathcal{S}^{tmp}(0)$ and gradually decrease the value of $\tau^{tmp}$ at $\mathcal{S}^{tmp}(t)$ to zero at end to achieve fully template internalization. Note that such compression system is also flexible, allowing it to halt at a desired non-zero compression rate. This flexibility allows to maintain a certain level of compressed template, serving as a trade-off to preserve inference accuracy in specific scenarios, as discussed in Section 5.3. In addition of the progressively decreasing template schedule, we also specify the template compressor $C$ for better utilization. we categorize it into two types which exactly reflects the primary components of the template defined in the problem formulation: the instruction compressor and document compressor:
1) Instruction Compressor targets the static elements within prompts, specifically focusing on the instructional content. Instructions in training data often consist of repeated directives, guidelines, or predefined tasks which are common across multiple training scenarios. The primary goal of the instruction compressor is to distill these instructions down to their essential components, eliminating verbosity and redundancy without compromising the clarity or intent of the instructions.
2) Document Compressor is designed to handle the bulkier and more detailed portions of the prompts, such as API documentation or static demonstrations. These sections typically include extensive technical descriptions and examples that, while informative, often contain a significant amount of repetitive or non-essential information Xu et al. (2023). The goal of the document compressor is to reduce the information unnecessary for understanding and applying the technical content, thereby streamlining the training process.
@@ 4.2 Example Absorption
Incorporating few-shot examples into fine-tuning not only improves information retrieval and memory recall Hübotter et al. (2024) but also yields substantial benefits in handling a variety of tasks with minimal data input Mosbach et al. (2023); Snell et al. (2017). However, directly adding lengthy few-shot examples to input prompts burdens the context window and increases inference latency. Motivated by this, we propose example absorption to benefit from the enhanced performance afforded by few-shot examples while prevent incurring significant additional overhead. Specifically, the example absorption mainly contains two stages: example retrieval and example removal.
1) Example Retrieval is designed to identify and select the most related few-shot examples from the training dataset and incorporate them into each training instance. The underlying rationale is to choose examples that closely align with the training instance so as to accelerate model’s internalization during training. We employ a straightforward approach that utilizes a relevance scoring function $s(\cdot,\cdot)$ to assess the similarity between examples and the training instance. Specifically, we select the top $k$ examples, varying at different training iterations, with the highest relevance scores to serve as our few-shot examples. For a training instance $(x_{i},y_{i})$ with $x_{i}$ being the input prompt and $y_{i}$ being the corresponding groundtruth output, the selection process can be expressed as follows:
@@
2) Example Removal aims to progressively internalize the prompt knowledge from few-shot examples into model parameters. To achieve this, we also adopt a predetermined schedule $\mathcal{S}^{egs}(t)$ to gradually decrease the number of demonstration examples in each prompt instance during the t-th iteration. Specifically, for a total of $T$ training iterations, we initially set $k$ examples at $\mathcal{S}^{egs}(0)$ and then gradually decrease the value of $k$ at each $\mathcal{S}^{egs}(t)$ to zero at end in order to achieve fully example internalization.
@@ 4.3 PromptIntern Pipeline
1) Preprocess. We first preprocess the input prompts to prepare them for the progressive training stage. Specifically, we process the prompt template to different compression rates based on the schedule $\mathcal{S}^{tmp}(t)$ and retrieve examples for each training instance based on the schedule $\mathcal{S}^{egs}(t)$. For better illustration, we provide an example of a pre-processed prompt with respect to schedule in Appendix B.
@@
Table 1: Comparison with Prompt Compression Approaches. The top-2 performance are highlighted with blue.
@@
2) Progressive Fine-tuning. We then fine-tune the model parameters for internalizing. Given the training iteration $t$, we update the model parameters as follows:
@@
3) Inference. After the progressive fine-tuning, we have trained the LLMs with updated model parameters $\theta_{T}$ to perform inference without adding instructions or any examples. Thus, we can predict the output simply with $f_{\theta_{T}}(x^{que})$.
@@ 5 Experiment
In this section, we evaluate the performance of PromptIntern across various benchmarks on the NL2Code task, which is widely recognized for its utility in evaluating LLMs on both fine-tuning efficacy and cost-effectiveness in real-world applications Zan et al. (2022). Our experiments primarily focus on two key perspectives: 1) Effectiveness: assessing the performance of textbf during inference phases; 2) Efficiency: quantifying the reduction in token usage and corresponding cost savings achievable through PromptIntern.
@@ 5.1 Settings
We use one-shot pass accuracy $Pass@1$ Austin et al. (2021) for MBPP, Exact Match (E.M.) for NL2F , and BLEU score for NL2Bash. In addition, we calculate the input token usage and compression rate $\tau$ for each dataset.
@@
To demonstrate the broad applicability of PromptIntern, we utilize both closed-source and open-source LLMs with different parameter sizes for fine-tuning and inference processes.1) Closed-Source: We apply GPT-4-0613 OpenAI (2023), abbreviated as GPT-4, and GPT-3.5-turbo-0125111https://platform.openai.com/docs/models/gpt-3-5-turbo, abbreviated as GPT-3.5. 2) Open-Source: We apply Mixtral-8x7B-v0.1 Jiang et al. (2024), abbreviated as Mixtral-8x7B, Llama2-7B Touvron et al. (2023), and Llama2-13B Touvron et al. (2023).
@@
Table 2: Comparison with direct fine-tuning baselines on MBPP datasets.
@@
Pass@1 Tokens Pass@1 Tokens Pass@1 Tokens Pass@1 Tokens Pass@1 Tokens
@@
Table 3: Comparison with direct fine-tuning baselines on NL2F datasets.
@@
E.M. Tokens E.M. Tokens E.M. Tokens E.M. Tokens E.M. Tokens
@@
Table 4: Comparison with direct fine-tuning baselines on NL2Bash datasets.
@@
BLEU Tokens BLEU Tokens BLEU Tokens BLEU Tokens BLEU Tokens
@@
5.2 Main results
Prompt Compression Approaches Comparison
Table 1 reports the overall result of PromptIntern with the prompt compression baselines inferenced on GPT-3.5 across all datasets. Here we establish the template compression rate $\tau_{tmp}$ at 0.3 across all prompt compression approaches as well as PromptIntern to ensure a fair comparison. And $\tau_{all}$ in the table represents the overall prompt’s compression rate. We observe that while utilizing a comparable number of tokens for inference, our approach significantly outperforms all baselines, achieving improvements of 5.6% on MBPP, 11.0% on NL2F, and 7.7% on NL2Bash. The result demonstrates that PromptIntern generally offers the best balance of efficiency and effectiveness across varied tasks. Note that since the Gist TokenMu et al. (2024) baseline is only applicable on open-source LLMs, we separately compare it with our approach which can be found at Appendix A.3.
Direct Fine-tuning Approaches Comparison
Table 2, 3, and 4 demonstrate the comparison results of PromptIntern with different direct fine-tuning baselines. For the MBPP dataset, our approach significantly outperforms No Template and Template baselines by 9.6%-10.8% and 0.6-1.3%, respectively, and achieve comparable performance against Template x0.6 with 2-shots baseline with far less token usage. In addition, our approach achieves a range of 9.8x-12.2x reduction in the number of input tokens required for comparable performance. Note that even though the template with 5-shots achieves the best performance, it requires 22.2x-27.4x more input tokens than our approach. This underscores the efficiency of our method in reducing the computational resources required for inference, while still delivering robust performance.
For the NL2F dataset, the results in Table 3 shows our approach greatly outperforms No Template baseline by 8.0%-19.0% with the same token usage for inference. In addition, our approach achieves comparable performance to the Template x0.6 with 2-shots baseline while reducing the required number of input tokens by 6.4x. Another finding from the result is that for LLMs with larger parameters, removing the prompt template causes less degradation in performance compared to smaller models, as seen in Template Only and No Template columns (-5.7% with LLama2-13B and -21.2% with LLama2-7B). This suggests that larger models have a better inherent capability to understand the input questions, even without detailed instructions provided by the prompt template.
For the NL2Bash dataset, we observe that under similar input tokens required, our approach also outperforms No Template baselines by 7.3%-11.3%, showing the superiority of prompt internalization over direct fine-tuning. Moreover, for reaching a similar performance as Template x0.6 with 5-shots, our approach reduces the required tokens for inference by around 15.5x. These results confirms the balance of cost efficiency and performance effectiveness of our PromptIntern approach during inference.
For additional experiments on inference speed of PromptIntern on each benchmark, please refer to Appendix A.4.

### v2

@@ Abstract
Recent advances in fine-tuning large language models (LLMs) have greatly enhanced their usage in domain-specific tasks. Despite the success, fine-tuning continues to rely on repeated and lengthy prompts, which escalate computational expenses, require more resources, and lead to slower inference. In this paper, we present a novel approach, PromptIntern, which internalizes prompt knowledge during model fine-tuning to achieve efficient inference and save costs. Instead of compressing the prompts for a vanilla model, PromptIntern aims to embed the recurrent prompt directly into the model parameters. We design a fine-tuning pipeline that includes instruction template compression, few-shot example absorption, and a progressive internalization strategy, effectively diminishing the need for intricate prompts during inference. Comprehensive experiments on challenging NL2Code tasks demonstrate that our method reduces input tokens by more than 90%, accelerates inference by 4.2 times, and reduces monetary inference costs by 88.3%.
PromptIntern: Saving Inference Costs by Internalizing Recurrent
Prompt during Large Language Model Fine-tuning
@@ 1 Introduction
Large language models (LLMs) have become pivotal in numerous natural language processing (NLP) applications, such as natural language generation Dong et al. (2019); Zheng et al. (2024), reasoning  Zhu et al. (2023); Sui et al. (2023), and code generation Luo et al. (2023b); He et al. (2024); Rozière et al. (2024). To enhance the predictive accuracy of LLMs in domain-specific tasks, recent techniques in fine-tuning, such as parameter-efficient fine-tuning (PEFT) He et al. (2021); Hu et al. (2021); Lester et al. (2021a), have been developed for pre-trained models to excel in specific tasks by adjusting their parameters to better align with targeted datasets Hu et al. (2021). Many of these fine-tuning approaches typically adopt prompts that are optimized and integrated with detailed instructions, examples, and retrieved documents through prompt engineering techniques such as chain-of-thought Wei et al. (2022), few-shot prompting Brown et al. (2020), and retrieval-augmented generation Lewis et al. (2020); Cheng et al. (2023).
@@
Although these advancements enhance the capabilities of LLMs during fine-tuning, they also present new challenges: Prompt engineering often necessitates longer prompts, and directly integrating lengthy prompts into the training process further increases computational costs during inference VM et al. (2024). This increase in cost precludes LLMs in many cost-sensitive scenarios where computational resources are constrained. Several prompt compression methods Li et al. (2023); Jiang et al. (2023a); Pan et al. (2024) have been proposed to reduce text redundancy. They design various prompt compression systems and strive to preserve maximum information between original and compressed prompts. While these methods ensure the retention of original prompt information, they mainly focus on the prompt perplexity and overlook the adaption of target LLMs during compression Pan et al. (2024). For challenging tasks that require model fine-tuning Mosbach et al. (2023), these approaches struggle to establish connections between compressed tokens and dynamically adjusted model parameters. Consequently, naively applying these compression methods often leads to large performance degradation, as relevant information may be inadvertently removed or distorted during the compression process.
In this paper, we propose a novel prompt internalization approach, namely PromptIntern, which internalizes prompt input during model fine-tuning and enables efficient inference. Unlike prompt compression which removes tokens based solely on prompt information entropy, we aim to transfer various types of prompt knowledge into updated model parameters, thereby directly enhancing LLMs’ understanding. Our idea is motivated by the human learning process as illustrated in Figure 1: During internship on-boarding, human interns need detailed instructions, examples, and documents to learn new tasks. As they internalize these materials and become familiar with their duties, they master the necessary skills and no longer require extra guidance. Similarly, when specific prompt information (e.g. task constraints, output formats/styles) is repeatedly exposed to an LLM during fine-tuning, the model can gradually internalize the knowledge into its updated parameters. Such repeated information can be progressively eliminated from prompt inputs since it becomes unnecessary for inference of the master LLM.
@@
  • We proposed PromptIntern [^1] Our code will be released at https://github.com/microsoft/PromptIntern, a novel prompt internalization method that aims to internalize repetitive prompt knowledge into the model’s parameters, achieving high inference efficiency while maintaining model performance.
  • We devised detailed prompt internalization strategies for template compression and example absorption along with a tailored progressive fine-tuning pipeline.
  • We conducted extensive experiments with detailed analyses on challenging NL2Code tasks. The experimental results show that our approach reduces token usage by over 90%, speeds up inference time by 4.2 times, and achieves 88.3% cost savings across a broad spectrum of LLMs.
@@ 2 Related Work
Prompt compression rephrases original prompts more concisely and is classified into task-aware and task-agnostic approaches. Specifically, the task-aware approaches, like LongLLMLingua Jiang et al. (2023b), utilize a question-aware coarse-to-fine-grained strategy to compress information based on the query. In contrast, methods like soft prompts Wingate et al. (2022); Liu et al. (2022); Mu et al. (2024) use learnable tokens to condense prompts. Conversely, the task-agnostic methods utilize metrics such as information entropy to eliminate redundant prompt information, with systems like LLMLingua Jiang et al. (2023a); Li et al. (2023) estimating token importance using a smaller model. Despite the demonstrated effectiveness of these methods, producing compressed text that can generalize across different tasks and be effectively integrated into training scenarios remains a challenge.
Model fine-tuning adopts pre-trained LLMs to specific tasks by modifying parameters. Based on the assumption that fine-tuning adds less new information to the model pre-trained on large internet-scale datasets, Parameter-Efficient Fine-Tuning (PEFT) methods aim to curtail the costs of tuning large models by adjusting a subset of parameters. Existing PEFT methods can be broadly categorized into three main approaches: 1) Adapter-based methods Houlsby et al. (2019); He et al. (2021): Introduce trainable modules within a static "backbone" network, offering flexibility but potentially increasing model size. 2) Prompt-based methods Lester et al. (2021b); Razdaibiedina et al. (2023); Nashid et al. (2023): Employ trainable "soft tokens" at input sequence start, requiring effective prompt design per task. 3) Low-rank adaptation methods Hu et al. (2021); Dettmers et al. (2024); Liu et al. (2024): Use low-rank matrices to approximate required weight adjustments, avoiding additional inference burden and often delivering strong performance. Despite advancements in fine-tuning strategies, data inputs should be carefully managed and distinguished from lengthy ones used for direct inference.
@@ 3 Problem Formulation
Let an input prompt as $x=(x^{tmp},x^{egs},x^{que})$, where each input prompt $x$ is considered as a tuple of three components: $x^{tmp}$ as the template such as fixed instructions, API docs, etc., $x^{egs}$ as the examples, and $x^{que}$ as the query. Typically, $x^{tmp}$ and $x^{egs}$ are relatively fixed and lengthy but essential for complex tasks. Let $f_{\theta}(\cdot)$ denote the neural network function of a LLM model, typically transformer Vaswani et al. (2017), parameterized by $\theta$. The generated output by LLM can be represented as $f_{\theta}(x)$.
@@ 4 Methodology
In this section, we introduce our method PromptIntern in detail. We first present the template compression to compress the entire fixed template part inside a prompt. Then we show the example absorption to effectively absorb demonstration examples into model parameters. Finally, we introduce a tailored training strategy for PromptIntern. The overall framework is shown in Figure 2.
@@ 4.1 Template Compression
We first introduce template compression, which is designed to compress the common template information exists across training instances. The motivation of the template compression stems from the following aspects: 1) Redundancy. The instruction is repetitive across prompts for a given task, often containing unnecessary tokens that do not contribute to the language model’s understanding, posing significant memory and computational burdens when the instruction is lengthy; and 2) Noise. Excessively long prompts may incorporate extraneous elements—either irrelevant or misleading information—that serve as noise and can adversely affect the model’s generation.
@@
where $C$ is a specific template compressor, $\tilde{x}^{tmp}$ is the compressed template, and $\tau^{tmp}$ is the template compression rate as defined in Jiang et al. (2023a), varying at differnt training interations. We then adopt a predetermined schedule $\mathcal{S}^{tmp}(t)$ to progressively reduce and internalize the prompt template information during the $t$-th training iteration. Specifically, for a total of $T$ training iterations, we initially set $\tau^{tmp}$ to 1 at $\mathcal{S}^{tmp}(0)$ and gradually decrease the value of $\tau^{tmp}$ at $\mathcal{S}^{tmp}(t)$ to zero at end to achieve fully template internalization. Note that such a compression system is also flexible, allowing it to halt at a desired non-zero compression rate. This flexibility allows to maintenance of a certain level of compressed template, serving as a trade-off to preserve inference accuracy in specific scenarios, as discussed in Section 5.4. In addition to the progressively decreasing template schedule, we also specify the template compressor $C$ for better utilization. we categorize it into two types which exactly reflect the primary components of the template defined in the problem formulation: the instruction compressor and document compressor:
Instruction Compressor targets the static elements within prompts, specifically focusing on the instructional content. Instructions in training data often consist of repeated directives, guidelines, or predefined tasks which are common across multiple training scenarios. The primary goal of the instruction compressor is to distill these instructions down to their essential components, eliminating verbosity and redundancy without compromising the clarity or intent of the instructions.
Document Compressor is designed to handle the bulkier and more detailed portions of the prompts, such as API documentation or static demonstrations. These sections typically include extensive technical descriptions and examples that, while informative, often contain a significant amount of repetitive or non-essential information Xu et al. (2023). The goal of the document compressor is to reduce the information unnecessary for understanding and applying the technical content, thereby streamlining the training process.
@@ 4.2 Example Absorption
Incorporating few-shot examples into fine-tuning not only improves information retrieval and memory recall Hübotter et al. (2024) but also yields substantial benefits in handling a variety of tasks with minimal data input Mosbach et al. (2023); Snell et al. (2017). However, directly adding lengthy few-shot examples to input prompts burdens the context window and increases inference latency. Motivated by this, we propose example absorption to benefit from the enhanced performance afforded by few-shot examples while preventing incurring significant additional overhead. Specifically, the example absorption mainly contains two stages: example retrieval and example removal.
Example Retrieval is designed to identify and select the most related few-shot examples from the training dataset and incorporate them into each training instance. The underlying rationale is to choose examples that closely align with the training instance so as to accelerate model’s internalization during training. We employ a straightforward approach that utilizes a relevance scoring function $s(\cdot,\cdot)$ to assess the similarity between examples and the training instance. Specifically, we select the top $k$ examples, varying at different training iterations, with the highest relevance scores to serve as our few-shot examples. For a training instance $(x_{i},y_{i})$ with $x_{i}$ being the input prompt and $y_{i}$ being the corresponding groundtruth output, the selection process can be expressed as follows:
@@
Example Removal aims to progressively internalize the prompt knowledge from few-shot examples into model parameters. To achieve this, we also adopt a predetermined schedule $\mathcal{S}^{egs}(t)$ to gradually decrease the number of demonstration examples in each prompt instance during the t-th iteration. Specifically, for a total of $T$ training iterations, we initially set $k$ examples at $\mathcal{S}^{egs}(0)$ and then gradually decrease the value of $k$ at each $\mathcal{S}^{egs}(t)$ to zero at end in order to achieve fully example internalization.
@@ 4.3 PromptIntern Pipeline
Preprocess. For the first step, We preprocess the input prompts to prepare them for the progressive training stage. Specifically, we process the prompt template to different compression rates based on the schedule $\mathcal{S}^{tmp}(t)$ and retrieve examples for each training instance based on the schedule $\mathcal{S}^{egs}(t)$. For better illustration, we provide an example of a pre-processed prompt with respect to schedule in Appendix C.
@@
Table 1: Comparison with prompt compression baselines on NL2Code benchmark. To ensure a fair comparison, we apply each baseline to compress the input prompt and use the compressed prompt as both training and testing data during model fine-tuning. We also standardize the compression ratio across all methods to approximately the same as indicated by $1/\tau_{all}$.
@@
(Inference on GPT-3.5) Pass@1 Input Tokens $1/\tau_{all}$ E.M. Input Tokens $1/\tau_{all}$ BLEU Input Tokens $1/\tau_{all}$
@@
Progressive Fine-tuning. We then fine-tune the model parameters for internalizing. Given the training iteration $t$, we update the model parameters as follows:
@@
Inference. After the progressive fine-tuning, we have trained the LLMs with updated model parameters $\theta_{T}$ to perform inference without adding instructions or any examples. Thus, we can predict the output simply with $f_{\theta_{T}}(x^{que})$.
@@ 5 Experiment
In this section, we evaluate the performance of PromptIntern across various benchmarks on the NL2Code task. The NL2Code task is widely recognized for its utility in evaluating LLMs on both fine-tuning efficacy and cost-effectiveness in real-world applications Zan et al. (2022). Following this, our experiments primarily focus on two key perspectives: 1) Effectiveness: assessing the performance accuracy of PromptIntern during inference phases; 2) Efficiency: quantifying the reduction in token usage and corresponding cost savings achievable through PromptIntern.
@@ 5.1 Settings
We use one-shot pass accuracy $Pass@1$ Austin et al. (2021) for MBPP, Exact Match (E.M.) for NL2F, and BLEU score for NL2Bash. We also calculate the input tokens usage and compression ratio $1/\tau$ for each dataset.
@@
To demonstrate the broad applicability of PromptIntern, we utilize both closed-source and open-source LLMs with different parameter sizes for fine-tuning and inference processes.1) Closed-Source: We apply GPT-4-0613 OpenAI (2023), abbreviated as GPT-4, and GPT-3.5-turbo-0125222https://platform.openai.com/docs/models/gpt-3-5-turbo, abbreviated as GPT-3.5. 2) Open-Source: We apply Mixtral-8x7B-v0.1 Jiang et al. (2024), abbreviated as Mixtral-8x7B, Llama2-7B Touvron et al. (2023), and Llama2-13B Touvron et al. (2023).
@@
Table 2: Comparison with direct fine-tuning baselines on NL2Code benchmark.
MBPP
@@
Pass@1 Input Tokens Pass@1 Input Tokens Pass@1 Input Tokens Pass@1 Input Tokens Pass@1 Input Tokens
@@
NL2F
@@
E.M. Input Tokens E.M. Input Tokens E.M. Input Tokens E.M. Input Tokens E.M. Input Tokens
@@
NL2Bash
@@
BLEU Input Tokens BLEU Input Tokens BLEU Input Tokens BLEU Input Tokens BLEU Input Tokens
@@
5.2 Prompt Compression Comparison
Table 1 reports the overall result of PromptIntern with the prompt compression baselines inferenced on GPT-3.5 across all datasets. Here we establish the template compression rate $\tau_{tmp}$ at 0.3 across all prompt compression approaches as well as PromptIntern to ensure a fair comparison. And $\tau_{all}$ in the table represents the overall prompt’s compression rate. We observe that while utilizing a comparable number of tokens for inference, our approach outperforms all baselines, achieving improvements of 5.6% on MBPP, 11.0% on NL2F, and 7.7% on NL2Bash. The result demonstrates that PromptIntern generally offers the best balance of efficiency and effectiveness across varied tasks. Note that since the Gist TokenMu et al. (2024) baseline is only applicable to open-source LLMs, we separately compare it with our approach which can be found in Appendix A.3.
5.3 Direct Fine-tuning Comparison
Table 2 shows the comparison of our approach with direct fine-tuning baselines on three datasets. In MBPP, our method outperforms the No Template and Template baselines by 9.6%-10.8% and 0.6%-1.3%, respectively, and achieves similar results to Template x0.6 with 2-shots, using fewer tokens. We also reduce input token usage by 9.8x-12.2x, compared to the Template with 5-shots, which requires 22.2x-27.4x more tokens. In NL2F, our method improves over No Template by 8.0%-19.0% with the same token usage, matching the Template x0.6 with 2-shots baseline while reducing token usage by 6.4x. We also observe that larger models (LLama2-13B) show less degradation without prompt templates compared to smaller models (LLama2-7B), with a difference of -5.7% verses -21.2%. In NL2Bash, our approach exceeds No Template by 7.3%-11.3% and reduces token usage by 15.5x compared to the baseline method Template x0.6 with 5-shots.